OpenStudy (anonymous):

Find the center and radius of the circle given by the equation: x^2+y^2+8x+10y+37=0. The center is ? The radius is ?

OpenStudy (anonymous):

\[x^{2}+8.x+y^{2}+10.y+37=0\] equivalent \[(x+4)^{2}-16 +(y+5)^{2}-25+32\] \[(x+4)^{2}+(y+5)^{2}=9\] C(-4,-5) and r=3

OpenStudy (anonymous):

37 instead 32,so the radius is 2

OpenStudy (moongazer):

when you are finding for the center of the radius and a circle. You need to identify first the equation if the center is at (h,k) or at (0,0). then you need to make them into standard form General forms: C (0,0) , x^2 + y^2 =r^2 C (h,k) , (x-h)^2 +(x-k)^2

OpenStudy (moongazer):

based on your question, the center is at (h,k) x^2+y^2+8x+10y+37=0 (x^2 + 8x) + (y^2 + 10y) = -37 (x^2 + 8x +16) +(y^2 +10y + 25) = -37 + 25 + 16 (x+4)^2 + (y+5)^2 = 4 C(-4,-5) r= 2 hope you understand it :)